How many grams of NiNi are formed from 56.6 g56.6 g of Ni2O3Ni2O3?

2Ni2O3(s)⟶4Ni(s)+3O2(g)2Ni2O3(s)⟶4Ni(s)+3O2(g)

Step 1: Show the strategy for solving this problem.

step 2: show the conversion for this problem and calculate grams of Ni.
Answer #1

Step 1:

strategy is:

mass of Ni2O3 —> mol of Ni2O3 —> mol of Ni —> mass of Ni

Step 2:

Molar mass of Ni2O3,

MM = 2*MM(Ni) + 3*MM(O)

= 2*58.69 + 3*16.0

= 165.38 g/mol

mass of Ni2O3 = 56.6 g

mol of Ni2O3 = (mass)/(molar mass)

= 56.6/1.654*10^2

= 0.3422 mol

Balanced chemical equation is:

2Ni2O3(s)⟶4Ni(s)+3O2(g)2Ni2O3(s)⟶4Ni(s)+3O2(g)

According to balanced equation

mol of Ni formed = (4/2)* moles of Ni2O3

= (4/2)*0.3422

= 0.6845 mol

Molar mass of Ni = 58.69 g/mol

mass of Ni = number of mol * molar mass

= 0.6845*58.69

= 40.17 g

Answer: 40.2 g